OpenStudy (anonymous):

Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements.

OpenStudy (anonymous):

@ElonaSushchik

OpenStudy (anonymous):

@chycora

OpenStudy (anonymous):

Gordon is evaluating some stocks he wants to purchase. You can see the function of the stock price below. The variable x represents days. Stock A B C Price f(x) = 25(1.08)x x f(x) 0 22.00 1 21.56 2 21.13 The initial price of this stock is $30, but it has been increasing 4% each day. Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements. Graph the price function for Stock D. Explain the key features of the graph with complete sentences. A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x). Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane. f(x + 2) f(–x) f(x) + 3

OpenStudy (anonymous):

Gordon is evaluating some stocks he wants to purchase. You can see the function of the stock price below. The variable x represents days. Stock A B C Price f(x) = 25(1.08)x x f(x) 0 22.00 1 21.56 2 21.13 The initial price of this stock is $30, but it has been increasing 4% each day. Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements. Graph the price function for Stock D. Explain the key features of the graph with complete sentences. A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x). Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane. f(x + 2) f(–x) f(x) + 3

OpenStudy (anonymous):

can u help

OpenStudy (elonasushchik):

okay soo Ok, to answer question (a), the price function of stock D, given that price is increasing by 4% each day, I am not sure about anything here, but maybe we can come out with something, is that exponential rise in price, or it is 0.4% of $30? If it is 0.4% 0f $30 , then h(x)=30+(0.04×30)x=30+1.2x

OpenStudy (elonasushchik):

OK, I believe it is a question on exponential. y(t)=a×ekt Where y(t) = value at time "t" a = value at the start k = rate of growth (when >0) or decay (when <0) t = time

OpenStudy (anonymous):

this right

OpenStudy (elonasushchik):

wayu mean yeh it's righ

OpenStudy (anonymous):

ok